# I-MCTS: Enhancing Agentic AutoML via Introspective Monte Carlo Tree Search

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have shown remarkable potential in automating machine learning tasks. However, existing LLM-based agents often struggle with low-diversity and suboptimal code generation. While recent work Chi et al. (2024) has introduced Monte Carlo Tree Search (MCTS) to address these issues, limitations persist in the quality and diversity of thoughts generated, as well as in the scalar value feedback mechanisms used for node selection. In this study, we introduce Introspective Monte Carlo Tree Search (I-MCTS), a novel approach that iteratively expands tree nodes through an introspective process that meticulously analyzes solutions and results from parent and sibling nodes. This facilitates a continuous refinement of the node in the search tree, thereby enhancing the overall decision-making process. Furthermore, we integrate a Large Language Model (LLM)-based value model to facilitate direct evaluation of each node’s solution prior to conducting comprehensive computational rollouts. A hybrid rewarding mechanism is implemented to seamlessly transition the Q-value from LLM-estimated scores to actual performance scores. This allows higher-quality nodes to be traversed earlier. Applied to the various ML tasks, our approach demonstrates a 6% absolute improvement in performance compared to the strong open-source AutoML agents, showcasing its effectiveness in enhancing agentic AutoML systems [^1] Resource available at https://github.com/jokieleung/I-MCTS.
@@ 1 Introduction
To address these challenges, we present Introspective Monte Carlo Tree Search (I-MCTS), a novel framework that introduces two key innovations. First, our introspective node expansion dynamically generates high-quality thought nodes through explicit analysis of parent and sibling node states. By incorporating reflective reasoning and feedback about prior solutions and their outcomes, this approach enables continuous quality improvement of the search tree nodes. Second, we develop a hybrid rewarding mechanism that combines: 1) LLM-estimated evaluations predicting node potential through a comprehensive machine learning evaluation criteria, and 2) actual performance scores on the development set from computational rollouts. Our adaptive reward blending strategy smoothly transitions Q-value from LLM-estimated values to actual values across search iterations. This facilitates higher-quality nodes to be traversed earlier.
Overall, the principal contributions of this work are threefold:
@@ 2.3 Hybrid Reward Mechanism
The primary objective of the evaluation phase is to assess the reward for the current node. For machine learning tasks, a conventional approach Schmidt et al. (2024); Chi et al. (2024) involves utilizing the performance metric on the development set as the reward. To derive a performance score of the given node involves rolling out from an intermediate state to a terminal state, which is quite costly since processes such as model training cost significant computational time. This limits the efficiency of node exploration. To address these challenges, we introduce a hybrid reward mechanism for node evaluation which combines LLM-estimated evaluations with actual performance scores to guide the search process more efficiently. For each node $x$, we employ an LLM-based value model $M_{\text{value}}$ to predict its potential performance. The value model takes as input the solution $\sigma_{\text{sol}}(x)$ and outputs an estimated score $s_{\text{LLM}}(x)$. This score is based on a comprehensive set of machine learning evaluation criteria, which is detailed in Appendix E.2. The LLM-estimated evaluation allows us to assess the quality of a node’s solution before conducting computationally expensive rollouts. This early evaluation helps prioritize nodes with higher potential, improving the efficiency of the search process. After the “simulation"" of a node $x$, we obtain the actual performance score $s_{\text{actual}}(x)$ based on the development set, which reflects the true performance of the node’s solution. To smoothly transition from LLM-estimated values to actual performance scores, we implement an adaptive reward blending strategy. The Q-value $Q(x)$ for a node $x$ is updated as follows:
@@ 2.4 Tree Search Process
Each rollout executes the full pipeline while caching intermediate results. The LLM value model provides real-time estimates $v_{\text{LLM}}$ before computational execution.
@@ 3 Experiment
We follow Chi et al. (2024) to benchmark our method on 20 machine learning datasets, including 13 classification tasks and 7 regression tasks from the AutoML Benchmark (AMLB)  Gijsbers et al. (2024) and Kaggle Competitions. All datasets are split into training, validation, and test sets with a 6:2:2 ratio, details refer to Appendix D. Each dataset is run three times using our method. We also use the Normalized Score (NS) defined in Appendix  C. We adopt Qwen2.5-72B-Instruct Yang et al. (2024) as our LLM. The temperature parameter is set to 0.2. $\gamma$ is set to 0.2, while $\alpha_{\text{explore}}$ is set to 2. Five baselines are included for comparison, including AutoGluon Tang et al. (2024), AutoSklearn Feurer et al. (2022), AIDE Schmidt et al. (2024), Data Interpreter Hong et al. (2024a), SELA Chi et al. (2024). We directly used the results reported in Chi et al. (2024). Experiments are conducted based on the MetaGPT Hong et al. (2024b) framework [^2] https://github.com/geekan/MetaGPT.
@@ 3.1 Main Results
The experimental results demonstrate the superior performance of the proposed I-MCTS approach compared to other strong AutoML baselines. As shown in Table 3, I-MCTS achieves the highest overall normalized score (NS) of 58.6% on average and 59.8% for the best predictions across 20 diverse tabular datasets. These results validate the effectiveness of I-MCTS’s unique introspective node expansion process, which enables continuous solution refinement through meticulous analysis of parent and sibling nodes. Also, the integration of an LLM-based value model and hybrid rewarding mechanism further contributes to the algorithm’s success by facilitating early identification and traversal of high-quality nodes. We include detailed results of each method in Appendix B.
@@ 3.2 Ablation Study
We follow Chi et al. (2024) to use boston, colleges, credit-g, Click_prediction_small, GesturePhaseSegmentationProcessed, and mfeat-factors to dataset for ablation study. We systematically evaluate the contribution of each component within our proposed Introspective Monte Carlo Tree Search (I-MCTS) framework. The results, as presented in Table 2, illustrate the incremental benefits of incorporating both the Introspective Node Expansion (INE) and the hybrid reward mechanism (HRM).
@@
Data Interpreter 56.4
SELA (Random Search) 58.6
SELA (MCTS) 60.9
I-MCTS (w/o INE) 61.1
I-MCTS (w/o HRM) 66.2
I-MCTS 66.8
Table 2: Ablation Study for each search setting on the selected 6 datasets. “I-MCTS (w/o INE)" means “without "Introspective Node Expansion", while “I-MCTS (w/o HRM)" means “without "hybrid reward mechanism".
@@ 3.3 Case Study
To demonstrate the superiority of our Introspective Monte Carlo Tree Search (I-MCTS) approach, we conduct a comprehensive visualization analysis of the tree search process using the kick dataset. As show in Appendix F, our empirical results reveal that I-MCTS effectively leverages the introspective information from previous nodes to generate task-specific and actionable insights, while SELA exhibit significant homogeneity and lack specificity. Furthermore, our hybrid reward mechanism enhances the exploration efficiency, enabling more effective identification of high-quality nodes within the same computational budget.
@@ Appendix C Evaluation Metrics
Appendix D Datasets
Table 4 outlines the detailed information of the datasets used for evaluation.
@@ Appendix D Datasets
Appendix E Prompts
E.1 Introspection Node Expansion Prompt
@@
E.2 Node Evaluation Prompt
@@
Appendix F Case Study
@@
F.1 Overview of I-MCTS’s search process
@@ 99 Simulated: False
F.2 Overview of baseline SELA’s search process

### v3

@@ Abstract
Recent advancements in large language models (LLMs) have shown remarkable potential in automating machine learning tasks. However, existing LLM-based agents often struggle with low-diversity and suboptimal code generation. While recent work Chi et al. (2024) has introduced Monte Carlo Tree Search (MCTS) to address these issues, limitations persist in the quality and diversity of thoughts generated, as well as in the scalar value feedback mechanisms used for node selection. In this study, we introduce Introspective Monte Carlo Tree Search (I-MCTS), a novel approach that iteratively expands tree nodes through an introspective process that meticulously analyzes solutions and results from parent and sibling nodes. This facilitates a continuous refinement of the node in the search tree. Furthermore, we integrate a Large Language Model (LLM)-based value model to facilitate direct evaluation of each node’s solution prior to conducting comprehensive computational rollouts. A hybrid rewarding mechanism is implemented to seamlessly transition the Q-value from LLM-estimated scores to actual performance scores. This allows higher-quality nodes to be traversed earlier. Applied to the various ML tasks, our approach demonstrates a 6% absolute improvement in performance compared to the strong open-source AutoML agents, showcasing its effectiveness in enhancing agentic AutoML systems [^1] Resource available at https://github.com/jokieleung/I-MCTS.
@@ 1 Introduction
In this paper, we present Introspective Monte Carlo Tree Search (I-MCTS), a novel inference-time scaling paradigm for Agentic AutoML that successfully combines the internal test-time compute (reflection) and external test-time compute (tree search) into one framework. There are two key innovations in our method. First, our introspective node expansion dynamically generates high-quality thought nodes through explicit analysis of parent and sibling node states. By incorporating reflective reasoning and feedback about prior solutions and their outcomes, this approach enables continuous quality improvement of the search tree nodes. Second, we develop a hybrid rewarding mechanism that combines: 1) LLM-estimated evaluations predicting node potential through a comprehensive machine learning evaluation criteria, and 2) actual performance scores on the development set from computational rollouts. Our adaptive reward blending strategy smoothly transitions Q-value from LLM-estimated values to actual values across search iterations. This facilitates higher-quality nodes to be traversed earlier. Overall, the principal contributions of this work are twofold:
@@ 2.3 Hybrid Reward Mechanism
The primary objective of the evaluation phase is to assess the reward for the current node. For machine learning tasks, a conventional approach Schmidt et al. (2024); Chi et al. (2024) involves utilizing the performance metric on the development set as the reward. To derive a performance score of the given node involves rolling out from an intermediate state to a terminal state, which is quite costly since processes such as model training cost significant computational time. This limits the efficiency of node exploration. To address these challenges, we introduce a hybrid reward mechanism for node evaluation which combines LLM-estimated evaluations with actual performance scores to guide the search process more efficiently. For each node $x$, we employ an LLM-based value model $M_{\text{value}}$ to predict its potential performance. The value model takes as input the solution $\sigma_{\text{sol}}(x)$ and outputs an estimated score $s_{\text{LLM}}(x)$. This score is based on a comprehensive set of machine learning evaluation criteria, which is detailed in Appendix H.2. The LLM-estimated evaluation allows us to assess the quality of a node’s solution before conducting computationally expensive rollouts. This early evaluation helps prioritize nodes with higher potential, improving the efficiency of the search process. After the "simulation" of a node $x$, we obtain the actual performance score $s_{\text{actual}}(x)$ based on the development set, which reflects the true performance of the node’s solution. To smoothly transition from LLM-estimated values to actual performance scores, we implement an adaptive reward blending strategy. The Q-value $Q(x)$ for a node $x$ is updated as follows:
@@ 2.4 Tree Search Process
Each rollout executes the full pipeline while caching intermediate results. The LLM value model provides real-time estimates $s_{\text{LLM}}$ before computational execution.
@@ 3 Experiment
For fair comparison, we follow Chi et al. (2024) to benchmark our method on 20 machine learning datasets, including 13 classification tasks and 7 regression tasks from the AutoML Benchmark (AMLB)  Gijsbers et al. (2024) and Kaggle Competitions. Five baselines are included for comparison, including AutoGluon Tang et al. (2024), AutoSklearn Feurer et al. (2022), AIDE Schmidt et al. (2024), Data Interpreter Hong et al. (2024a), SELA Chi et al. (2024). We directly used the results reported in Chi et al. (2024). Experiments are conducted based on the MetaGPT Hong et al. (2024b) framework [^2] https://github.com/geekan/MetaGPT. More experimental details please refer to the Appendix D.
@@ 3.1 Main Results
The experimental results demonstrate the superior performance of the proposed I-MCTS approach compared to other strong AutoML baselines. As shown in Table 3, I-MCTS achieves the highest overall normalized score (NS) of 58.6% on average and 59.8% for the best predictions across 20 diverse tabular datasets. These results validate the effectiveness of I-MCTS’s unique introspective node expansion process, which enables continuous solution refinement through meticulous analysis of parent and sibling nodes. We include detailed results of each method in Appendix B.
@@ 3.2 Ablation Study
Performing ablation on all datasets is very time-consuming, hence, we employ a subset of datasets to do an ablation study. We follow Chi et al. (2024) to choose the first two datasets alphabetically for each machine learning task type (Classification / Multi-class / Regression ). The results in Table 2 illustrate the incremental benefits of incorporating both the Introspective Node Expansion (INE) and the hybrid reward mechanism (HRM). Specifically, the results between I-MCTS (w/o HRM, rollout=5) and I-MCTS (rollout=5) shows that, incorporating an LLM-estimated evaluation score facilitates better solutions to be explored earlier, especially under a lower resource scenario.
@@
Data Interpreter (r=10) 56.4
SELA (Random Search, r=10) 58.6
SELA (MCTS, r=10) 60.9
I-MCTS (w/o INE, r=10) 61.1
I-MCTS (w/o HRM, r=10) 66.2
I-MCTS (r=10) 66.8
I-MCTS (w/o HRM, r=5) 61.9
I-MCTS (r=5) 64.7
Table 2: Ablation Study for each search setting on the selected 6 datasets. "w/o INE" means "without "Introspective Node Expansion", while "w/o HRM" means “without "hybrid reward mechanism". "r" means the rollout numbers.
@@ 3.3 Case Study
We conduct a comprehensive visualization analysis of the tree search process using the kick dataset. As show in Appendix I, our empirical results reveal that I-MCTS effectively leverages the introspective information from previous nodes to generate task-specific and actionable insights, while SELA exhibit significant homogeneity and lack specificity. Furthermore, our hybrid reward mechanism enhances the exploration efficiency, enabling more effective identification of high-quality nodes within the same computational budget.
@@ Appendix C Evaluation Metrics
Appendix D Experimental Details
All datasets are split into training, validation, and test sets with a 6:2:2 ratio. For the LLM-based Agents (I-MCTS, SELA, Data Interpreter, AIDE, etc), we employ a consistent initial task prompt across all methods  [^3] Details can be refer to our code repository https://anonymous.4open.science/r/I-MCTS. This prompt encompasses the dataset name, target column, and evaluation metric. We choose Qwen2.5-72B-Instruct as our foundation LLM, we set the temperature parameter to 0.5. $\alpha_{\text{explore}}$ is set to 2. AIDE conducts 10 iterations per execution, while I-MCTS, SELA perform 10 rollouts. For I-MCTS and SELA, we employ Data Interpreter as the experimenter. The value of $\gamma$ in HRM is set to 0.2 across all datasets. This value is to control the decay rate of the LLM estimated score. Larger $\gamma$ makes the influence of the LLM estimated to persist longer, while smaller $\gamma$ allows the actual performance results to dominate the hybrid reward score more quickly. $\alpha_{explore}$ is set to 2. This aims to balance exploration and exploitation in the method’s search strategy. Following Chi et al. (2024), each method, except for AutoGluon, is run three times for each dataset. AutoGluon, being deterministic, is run only once with its default settings. AutoSklearn is also run with default settings, limited to 600 seconds per task.
Appendix E Datasets
Table 4 outlines the detailed information of the datasets used for evaluation. For ablation study, we use boston, colleges, credit-g, Click_prediction_small, GesturePhaseSegmentationProcessed, and mfeat-factors.
@@ Appendix E Datasets
Appendix F Cost-efficiency analysis
Our ablation studies across six datasets demonstrate that with rollouts=10, our method incurs about 30% higher token consumption than SELA (averaging \$0.065 vs \$0.05 per dataset), which we consider comparatively equitable, given the empirical challenges in maintaining identical token consumption with baselines. More critically, we consider that experimental iteration counts constitute a more fundamental constraint than token budgets, as each agent trial necessitates substantial computational resource utilization (e.g., CPU/GPU allocation for model training). The relative significance of token costs will be anticipated to diminish with advancements in AI infrastructure and architectural optimizations.
Appendix G Related work
Recent advances in autonomous language agents have focused on iterative self-improvement and multi-agent coordination. Madaan et al. (2023) proposed SELF-REFINE, which iteratively refines LLM outputs through self-generated feedback without additional training. Building on iterative refinement, Shinn et al. (2023) introduced Reflexion, which leverages natural language feedback and episodic memory to enable LLM agents to learn from errors. Zhou et al. (2024) unified reasoning and action through Language Agent Tree Search (LATS). This is the most related work that also introduces self-reflection techniques into the Monte Carlo Tree Search process. However, the reflection is generated and used after the trajectory fails as additional context for future trials. By contrast, our introspective node expansion dynamically generates high-quality thought nodes through explicit analysis of parent and sibling node states, which enables continuous quality improvement of the search tree nodes.
Appendix H Prompts
H.1 Introspection Node Expansion Prompt
@@
H.2 Node Evaluation Prompt
@@
Appendix I Case Study
@@
I.1 Overview of I-MCTS’s search process
@@ 99 Simulated: False
I.2 Overview of baseline SELA’s search process